PURE ROTATIONAL SPECTRA OF THE REACTION PRODUCTS OF LASER ABLATED THORIUM METAL AND OXYGEN MOLECULES ENTRAINED WITHIN SUPERSONIC EXPANSIONS OF NOBLE GASES

\underline{B. E. LONG}, Department of Chemistry, Wesleyan University, 52 Lawn Avenue, Middletown, Connecticut, 06459; S. A. COOKE, School of Natural and Social Sciences, Purchase College SUNY, 735 Anderson Hill Road, Purchase, NY 10577, USA.

Thorium metal has been laser ablated in the presence of pure oxygen entrained in high pressure argon. The products of the ablation event have been supersonically expanded into the Fabry-Pérot cavity of a time-domain microwave spectrometer. New measurements have been performed on thorium monoxide, where the J = 1 ← 0 transition has been recorded in the v = 8, 9, and 10 vibrational levels. Further to this, a thorium-dependent spectral transition has been observed at 19251.8740(10) MHz. The carrier of this signal is to be determined, however, it requires an oxygen concentration of 0.1 % which is significantly higher than the optimal oxygen concentration required for observation of the thorium monoxide transitions. It is postulated that the presently unidentified thorium-containing compound is ThO_{2} which has C_{2}v symmetry. Progress on the measurement of this spectra, together with other thorium experiments and supporting quantum chemical calculations, will be presented.